# DAO: A New and Public Non-Relativistic Reflection Model

### v1

@@ 2.4 Equilibrium and iteration
Considering that XSTAR is the most time-consuming component of the calculation and that we employ a high energy resolution, we do not use the Accelerated Lambda Iteration (ALI) method in this work. In our case, the accelerated $\Lambda^{*}$-operator becomes a full matrix of dimension $N_{\tau_{T}}\times N_{E}\times N_{E}$ because the partial redistribution function (specifically, the scattering redistribution function in our context) is considered (1995A&A...297..771P), making the corresponding linear system prohibitively expensive to solve. Moreover, 2012A&A...545A.120S demonstrated that ALI still requires a large number of iterations when the exact Compton scattering redistribution function is used.
@@
3 results
3.1 Comparison with Other Models
First, we employ the nthcomp model (1996MNRAS.283..193Z; 1999MNRAS.309..561Z) as the illuminating source at the top boundary. We set $I_{\mathrm{type}}=\mathrm{file}$ and obtain the incident spectrum by setting the reflection fraction to 0, $\Gamma=2.0$, and $kT_{E}=60$ keV in xillvercp; this spectrum is then adopted as the standard incident radiation field in this paper. The results are compared with the reflionx_hd_nthcomp (2007MNRAS.381.1697R; 2020MNRAS.498.3888J) and xillvercp tables. To match these models, we do not adopt any bottom incident radiation field; thus, we set $B_{\mathrm{sw}}=0$. Our model and xillvercp adopt the solar abundances table from 1996ASPC...99..117G for elements from H to Zn, in contrast to the reflionx model, which uses the abundances from 1983ApJ...270..119M. Accordingly, in this section, we set $A_{\mathrm{Fe}}$ in our model and xillvercp to 1.32 to match the setting in reflionx. In addition, we set $\log\xi=1.0$ (top row), 2.0 (middle row), and 3.0 (bottom row), and adopt both the Gaussian approximation (left column) and the exact energy redistribution function (right column) for Compton scattering in Figure 1. All other parameters are set to their default values listed in Table 1.
The spectra in Figure 1 are plotted in the form of $EJ_{E}$. The reflionx model assumes an isotropic radiation field, whereas both DAO and xillver compute an anisotropic radiation field. For DAO and xillver, the mean intensity is evaluated as $J(E)=\int I(\mu,E)\,d\mu$, where $\mu$ is sampled according to the internal angular discretization adopted in each model [^8] If we choose $N_{\mu}=10$, our angle grid is the same as xillver’s: $\mu=\{0.05,0.15,0.25,0.35,0.45,0.55,0.65,0.75,0.85,0.95\}$.. As illustrated in Figure 1, the outputs of the three models are consistent within the 0.1 to 1000 keV range. All three models exhibit a prominent Compton hump, strong absorption edges and emission lines. In particular, in the spectrum generated using the exact redistribution, the Compton hump drops more rapidly at higher energies. In Section 3.2, we analyze in detail the impact of different redistribution functions.
@@ 3.1 Comparison with Other Models
DAO ATDB (2024-10-17T14:09:13)(2021Atoms...9...12M) 2.59(2021ApJ...908...94K)
xillvercp ATDB (2012-08-03T01:58:54) 2.2.1bn (2001ApJS..133..221K)
reflionx custom definition (see 2005MNRAS.358..211R) not used
@@
We also test our model with a softer and a harder illumination, specifically $\Gamma=1.4$ and $\Gamma=2.4$, with the results presented in Figure 3. In Figure 3, the ionization parameter is set at 3.0 for all spectra; the other parameters are the same as in Figure 1. Hard X-ray illumination with exact scattering redistribution displays a notable difference in the Compton hump from that of the Gaussian-approximated redistribution. This is caused by the inaccuracy of the Gaussian-approximation redistribution function for high-energy photons and electrons (see Section 3.2 for more details). However, we also find noticeable discrepancies among the three models in the soft X-ray band (0.1–1 keV) for any nthcomp incident spectrum. Light-element $\alpha$ lines (e.g., C, N, O) as well as Fe L-shell emission lines are generally present in this energy range. For reflionx, such discrepancies are expected because the model includes only a limited set of ions. In contrast, the differences in the soft X-ray range between xillver and DAO most likely arise from the use of different atomic databases and different versions of XSTAR in their calculations. Table 3 summarizes the XSTAR versions and atomic data tables employed by each code. The ATDB 2024 release incorporates corrections for high-density plasma effects and improved radiative and collisional rates for odd-$Z$ elements, with related developments reported in 2017A&A...604A..63M; 2018A&A...616A..62M; 2021Atoms...9...12M. Moreover, 2024ApJ...974..280D demonstrated the impact of the latest XSTAR and ATDB versions on reflection spectra, as shown in their Figures 3 and 4. They also found that using different versions of XSTAR and atomic databases in the xillver model produced noticeable differences.
@@
In summary, DAO provides a more accurate treatment of atomic processes by employing the latest versions of XSTAR and ATDB, and it achieves a more precise description of Compton scattering through the use of the exact redistribution function. Furthermore, we compare DAO with pexrav (1995MNRAS.273..837M) to validate our treatment of Compton scattering. The results are shown in Figure 17 in Appendix A. Although the pexrav spectrum exhibits a slightly higher peak in the Compton hump, the two models are consistent at energies $E<20$ keV and $E>50$ keV.
3.2 Effects of Energy Redistribution Functions
@@
Compton scattering is a stochastic process that alters both the propagation direction and the energy of photons, and it should be described using the exact quantum mechanical formalism. However, models such as xillver and reflionx adopt a Gaussian approximation (1978ApJ...219..292R) to treat this process, which is only valid to describe the energy redistribution for low-energy photons ($E_{\rm ph}\ll m_{e}c^{2}$) scattered by low-temperature electron clouds ($kT_{e}\ll E_{\rm ph}$). Appendix Figure 18 illustrates the differences between the exact and Gaussian redistribution functions. (1) At low electron temperatures and for soft incident photons ($E_{i}=1$ keV), both the Gaussian and exact redistribution functions exhibit a single peak with good symmetry. (2) When the incident photon energy approaches $m_{e}c^{2}$ while the electron temperature remains low, the exact redistribution function develops a double-peaked and asymmetric structure. (3) At sufficiently high electron temperatures (e.g., $T=10^{9}$ K), the Gaussian approximation becomes excessively broadened and no longer captures the actual shape of the Compton scattering redistribution. 2020ApJ...897...67G investigated this effect on the reflection spectrum, but it has not been implemented in a publicly released table model before.
@@
The exact redistribution function for Compton scattering has been extensively studied. 1981MNRAS.197..451G incorporated the relativistic Maxwellian velocity distribution into the redistribution function, but a computational error was present in the original paper. Subsequently, 1993AstL...19..262N, 1996ApJ...470..249P, and 2010ApJS..189..286P investigated the redistribution function for photons scattered by anisotropic electrons using quantum electrodynamical methods. Later, 2017MNRAS.469.2032M corrected the earlier error in 1981MNRAS.197..451G and demonstrated consistency among these methods. 2020ApJ...897...67G used the method described in 1993AstL...19..262N and showed reflection spectra using the exact redistribution function. We present the algorithms for the Gaussian-approximated and quantum-mechanical redistribution functions in Appendix B.
@@
Current instruments, such as NuSTAR, can image the sky in the 3–79 keV band, while Insight-HXMT extends observations to hard X-ray energies up to 250 keV (2020SCPMA..6349502Z). Both of them can observe the hard X-ray band, where deviations in measurements arise, especially for hard spectra and high coronal temperatures (2017MNRAS.468.3489K; 2019MNRAS.490.1350B). We therefore employ the exact redistribution function for all subsequent calculations.
3.3 Effects of varying the ionization parameter
@@
We assume a constant-density atmosphere, and the form of 1969ApJ...156..943T is used to calculate the ionization parameter $\xi$ (Eq. 2). At each optical depth layer, the total ionizing flux $F_{x}$ of the radiation field is calculated to obtain the ionization parameter. Figure 5 illustrates the emergent spectra for various ionization parameters, considering the exact redistribution function for Compton scattering. No normalization has been applied to the final spectra displayed in this plot. It is evident that spectra corresponding to a higher ionization parameter exhibit greater luminosity, whereas spectra with a lower ionization parameter display more prominent atomic line features and the Compton shoulder.
Figure 6 shows the temperature profile as a function of the Thomson optical depth (from 10-4 to 10) for the two incident spectra shown in Figure 5. In a highly ionized atmosphere, the dominant heating process is electron recoil following Compton scattering. The net heating rate is given by 1978ApJ...219..292R
@@
3.4 Effects of varying the incident angle
In our model, the coronal illumination field is specified by a single incidence angle. The $I_{\mathrm{inc}}$ at the top boundary of the second-order radiative transfer equation (Eq. 12) is set to zero everywhere except at $\mu_{\mathrm{inc}}$. Since the flux is defined as the first moment of the intensity,
@@ 3.4 Effects of varying the incident angle
Figure 10 shows the emission-angle-averaged flux for different incidence angles. Incidence angles are sampled from 0.1 to 0.9 in steps of 0.1 (color-coded from purple to yellow). The illumination radiation field is modeled with a cut-off power-law, using $\Gamma=2.0$ and $E_{\mathrm{cut}}=300~\mathrm{keV}$. The ionization parameter is set to 2.0, and the number of angular grids is fixed at $N_{\mu}=30$ to improve the angular resolution, while all other model parameters are set to their default values. The results are broadly consistent with our expectation: the gas temperature and the strength of the radiation hump show a clear inverse correlation with $\cos\theta_{\mathrm{inc}}$.
@@
3.5 Emergent intensity along all angle bins
Depending on the choice of angular binning, the emergent intensity can be obtained for different emission angles. In the special case where $\mu_{\mathrm{inc}}=\mu_{\mathrm{emis}}$, the emergent radiation field is given by
@@ 3.5 Emergent intensity along all angle bins
As shown in Figure 12, the radiation intensity decreases as $\mu$ increases. This behavior is primarily driven by the effective optical depth, defined as
@@
Photons escaping at grazing angles (low $\mu$) traverse a longer effective path through the medium. Consequently, spectral features originating from deeper layers, such as emission lines from heavy elements (e.g., Fe), suffer stronger attenuation compared to those produced near the surface (e.g., C, O). For hard X-ray photons, the extended path length increases the probability of scattering, rendering the Compton hump highly sensitive to minor variations in optical depth. This effect, well established in previous works (1981ApJ...248..738L; 1988ApJ...335...57L; 1995MNRAS.273..837M), is critical for interpreting the continuum. Indeed, Figure 12 reveals that the Compton hump is most pronounced at the smallest cosine of the emergent angle.
The angular dependence of the emissivity plays a vital role in determining the observed spectrum (e.g., 2025MNRAS.536.2594L). Recently, 2025ApJ...989..168H highlighted the impact of emission angles on relativistic spectra shaped by strong gravity. In Appendix C, we present relativistic spectra calculated using the formalism adopted from the relconv, relxill and relxillA model.
@@
3.6 Effects of varying the hydrogen density
Currently, DAO assumes reflection in a constant-density slab, with a default hydrogen density of 1015 cm-3. In AGN or XRB systems, the disk densities are sometimes greater than 1015 cm-3 (2019MNRAS.489.3436J) and sometimes even higher than 1020 cm-3 (2023ApJ...951..145L). However, the standard XSTAR atomic table (ADTB) only supports hydrogen densities lower than 1018 cm-3; a specific atomic database is needed for higher hydrogen densities (2021ApJ...908...94K). DAO will be updated to consider higher densities once the atomic database is available.
In Figure 13, we compare reflection spectra for hydrogen densities of 1015, 1016, and 1017 cm-3 under nthcomp illumination. Other parameters are set at their default values. The results are consistent with the findings of 2016MNRAS.462..751G: for a given ionization parameter, the gas temperature increases with hydrogen density, and the continuum emission in the low energy band displays clear differences among these spectra.
To further study the density effects in the soft X-ray band and temperature profile, we perform additional calculations with a blackbody incident spectrum of temperature 1 keV, which allows us to focus on the soft X-ray band. As shown in Figure 14, the Compton hump is weaker under this type of illumination than for a power-law incident spectrum, and the high-energy parts of the spectra become nearly identical across different densities. However, noticeable differences persist in the low-energy continuum emission. The temperature profiles show obvious differences at large optical depths.

### v2

@@ II.4 Equilibrium and iteration
Considering that XSTAR is the most time-consuming component of the calculation and that we employ a high energy resolution, we do not use the Accelerated Lambda Iteration (ALI) method in this work. In our case, the accelerated $\Lambda^{*}$-operator becomes a full matrix of dimension $N_{\tau_{T}}\times N_{E}\times N_{E}$ because the partial redistribution function (specifically, the scattering redistribution function in our context) is considered (Paletou and Auer 1995), making the corresponding linear system prohibitively expensive to solve. Moreover, Suleimanov et al. (2012) demonstrated that ALI still requires a large number of iterations when the exact Compton scattering redistribution function is used
@@
II.5 Implementation
In summary, the DAO code utilizes the public XSTAR v2.59 (Kallman and Bautista 2001; Kallman et al. 2021) to treat atomic physics. The differences in the treatment of atomic physics among DAO, xillver, and reflionx are summarized in Table 3. For coronal illumination, the cut-off power law, power law, blackbody, and custom-defined spectra are implemented. In addition, we incorporate public code from XSPEC (Grevesse et al. 1996) nthcomp (Zdziarski et al. 1996; Życki et al. 1999) and comptt (Titarchuk 1994) as illumination spectra. For Compton scattering, we use either a pre-calculated exact or Gaussian-approximated redistribution function. The exact redistribution function is calculated using the GitHub repository provided by García et al. (2020). For radiative transfer, we employ the FTM, which has been widely applied in the scenario of X-ray reflection (Nayakshin et al. 2000; García and Kallman 2010; García et al. 2013). Due to the proprietary nature of the xillver code, we can only compare our code with its public table models, which will be discussed in the next section.
@@ II.5 Implementation
III results
III.1 Comparison with Other Models
First, we employ the nthcomp model (Zdziarski et al. 1996; Życki et al. 1999) as the illuminating source at the top boundary. We set $I_{\mathrm{type}}=\mathrm{file}$ and obtain the incident spectrum by setting the reflection fraction to 0, $\Gamma=2.0$, and $kT_{E}=60$ keV in xillvercp; this spectrum is then adopted as the standard incident radiation field in this paper. The results are compared with the reflionx_hd_nthcomp (Ross and Fabian 2007; Jiang et al. 2020) and xillvercp tables. To match these models, we do not adopt any bottom incident radiation field; thus, we set $B_{\mathrm{sw}}=0$. Our model and xillvercp adopt the solar abundances table from Grevesse et al. (1996) for elements from H to Zn, in contrast to the reflionx model, which uses the abundances from Morrison and McCammon (1983). Accordingly, in this section, we set $A_{\mathrm{Fe}}$ in our model and xillvercp to 1.32 to match the setting in reflionx. In addition, we set $\log\xi=1.0$ (top row), 2.0 (middle row), and 3.0 (bottom row), and adopt both the Gaussian approximation (left column) and the exact energy redistribution function (right column) for Compton scattering in Figure 1. All other parameters are set to their default values listed in Table 1.
The spectra in Figure 1 are plotted in the form of $EJ_{E}$. The reflionx model assumes an isotropic radiation field, whereas both DAO and xillver compute an anisotropic radiation field. For DAO and xillver, the mean intensity is evaluated as $J(E)=\int I(\mu,E)\,d\mu$, where $\mu$ is sampled according to the internal angular discretization adopted in each model [^8] If we choose $N_{\mu}=10$, our angle grid is the same as xillver’s: $\mu=\{0.05,0.15,0.25,0.35,0.45,0.55,0.65,0.75,0.85,0.95\}$.. As illustrated in Figure 1, the outputs of the three models are consistent within the 0.1 to 1000 keV range. All three models exhibit a prominent Compton hump, strong absorption edges and emission lines. In particular, in the spectrum generated using the exact redistribution, the Compton hump drops more rapidly at higher energies. In Section III.2, we analyze in detail the impact of different redistribution functions. In principle, for the spectra calculated by Gaussian-approximated redistribution function, our results should be identical to xillverCp since we employ the same mathematical formalism, but in practice the different versions of XSTAR employed and other small differences in assumptions are enough to cause small discrepancies in the outputs of the two models. Still, as expected, DAO agrees with xillver better than with reflionx.
@@ III.1 Comparison with Other Models
DAO ATDB (2024-10-17T14:09:13)(Mendoza et al. 2021) 2.59(Kallman et al. 2021)
xillvercp ATDB (2012-08-03T01:58:54) 2.2.1bn (Kallman and Bautista 2001)
reflionx custom definition (see Ross and Fabian 2005) not used
@@
We also test our model with a softer and a harder illumination, specifically $\Gamma=1.4$ and $\Gamma=2.4$, with the results presented in Figure 3. In Figure 3, the ionization parameter is set at 3.0 for all spectra; the other parameters are the same as in Figure 1. Hard X-ray illumination with exact scattering redistribution displays a notable difference in the Compton hump from that of the Gaussian-approximated redistribution. This is caused by the inaccuracy of the Gaussian-approximation redistribution function for high-energy photons and electrons (see Section III.2 for more details). However, we also find noticeable discrepancies among the three models in the soft X-ray band (0.1–1 keV) for any nthcomp incident spectrum. Light-element $\alpha$ lines (e.g., C, N, O) as well as Fe L-shell emission lines are generally present in this energy range. For reflionx, such discrepancies are expected because the model includes only a limited set of ions. In contrast, the differences in the soft X-ray range between xillver and DAO most likely arise from the use of different atomic databases and different versions of XSTAR in their calculations. Table 3 summarizes the XSTAR versions and atomic data tables employed by each code. The ATDB 2024 release incorporates corrections for high-density plasma effects and improved radiative and collisional rates for odd-$Z$ elements, with related developments reported in Mendoza et al. (2017, 2018, 2021). Moreover, Ding et al. (2024) demonstrated the impact of the latest XSTAR and ATDB versions on reflection spectra, as shown in their Figures 3 and 4. They also found that using different versions of XSTAR and atomic databases in the xillver model produced noticeable differences.
@@
In summary, DAO incorporates the more accurate treatment of atomic processes originally explored in Ding et al. (2024) by employing the latest versions of XSTAR, and the precise description of Compton scattering first explored in García et al. (2020) into a publicly available code that can be run in bespoke configurations. Furthermore, we compare DAO with pexrav (Magdziarz and Zdziarski 1995) to validate our treatment of Compton scattering. The results are shown in Figure 17 in Appendix A. Although the pexrav spectrum exhibits a slightly higher peak in the Compton hump, the two models are consistent at energies $E<20$ keV and $E>50$ keV.
III.2 Effects of Energy Redistribution Functions
Compton scattering is a stochastic process that alters both the propagation direction and the energy of photons, and it should be described using the exact quantum mechanical formalism.
@@ III.2 Effects of Energy Redistribution Functions
However, all publicly available reflection table models adopt a Gaussian approximation (Ross et al. 1978) to treat this process, which is only valid to describe the energy redistribution for low-energy photons ($E_{\rm ph}\ll m_{e}c^{2}$) scattered by low-temperature electron clouds ($kT_{e}\ll E_{\rm ph}$). Appendix Figure 18 illustrates the differences between the exact and Gaussian redistribution functions. (1) At low electron temperatures and for soft incident photons ($E_{i}=1$ keV), both the Gaussian and exact redistribution functions exhibit a single peak with good symmetry. (2) When the incident photon energy approaches $m_{e}c^{2}$ while the electron temperature remains low, the exact redistribution function develops a double-peaked and asymmetric structure. (3) At sufficiently high electron temperatures (e.g., $T=10^{9}$ K), the Gaussian approximation becomes excessively broadened and no longer captures the actual shape of the Compton scattering redistribution. García et al. (2020) investigated this effect on the reflection spectrum and included it into the proprietary xillver source code, but it has not yet been implemented in a publicly released table model. DAO incorporates the publicly available source code released by García et al. (2020) to include the exact Compton scattering kernel.
@@
The exact redistribution function for Compton scattering has been extensively studied. Guilbert (1981) incorporated the relativistic Maxwellian velocity distribution into the redistribution function, but a computational error was present in the original paper. Subsequently, Nagirner and Poutanen (1993), Poutanen and Svensson (1996), and Poutanen and Vurm (2010) investigated the redistribution function for photons scattered by anisotropic electrons using quantum electrodynamical methods. Later, Madej et al. (2017) corrected the earlier error in Guilbert (1981) and demonstrated consistency among these methods. García et al. (2020) used the method described in Nagirner and Poutanen (1993) and showed reflection spectra using the exact redistribution function. We present the algorithms for the Gaussian-approximated and quantum-mechanical redistribution functions in Appendix B.
[…]
@@
Current instruments, such as NuSTAR, can image the sky in the 3–79 keV band, while Insight-HXMT extends observations to hard X-ray energies up to 250 keV (Zhang et al. 2020). Both of them can observe the hard X-ray band, where deviations in measurements arise, especially for hard spectra and high coronal temperatures (Kara et al. 2017; Buisson et al. 2019). We therefore employ the exact redistribution function for all subsequent calculations.
III.3 Effects of varying the ionization parameter
@@
We assume a constant-density atmosphere, and the form of Tarter et al. (1969) is used to calculate the ionization parameter $\xi$ (Eq. 2). At each optical depth layer, the total ionizing flux $F_{x}$ of the radiation field is calculated to obtain the ionization parameter. Figure 5 illustrates the emergent spectra for various ionization parameters, considering the exact redistribution function for Compton scattering. No normalization has been applied to the final spectra displayed in this plot. It is evident that spectra corresponding to a higher ionization parameter exhibit greater luminosity, whereas spectra with a lower ionization parameter display more prominent atomic line features and the Compton shoulder.
Figure 6 shows the temperature profile as a function of the Thomson optical depth (from 10-4 to 10) for the two incident spectra shown in Figure 5. In a highly ionized atmosphere, the dominant heating process is electron recoil following Compton scattering. The net heating rate is given by Ross et al. (1978)
@@
III.4 Effects of varying the incident angle
In our model, following the xillver models, the coronal illumination field is specified by a single incidence angle. The $I_{\mathrm{inc}}$ at the top boundary of the second-order radiative transfer equation (Eq. 12) is set to zero everywhere except at $\mu_{\mathrm{inc}}$. Since the flux is defined as the first moment of the intensity,
@@ III.4 Effects of varying the incident angle
Figure 10 shows the emission-angle-averaged flux for different incidence angles. Incidence angles are sampled from 0.1 to 0.9 in steps of 0.1 (color-coded from purple to yellow). The illumination radiation field is modeled with a cut-off power-law, using $\Gamma=2.0$ and $E_{\mathrm{cut}}=300~\mathrm{keV}$. The ionization parameter is set to 2.0, and the number of angular grids is fixed at $N_{\mu}=30$ to improve the angular resolution, while all other model parameters are set to their default values. The results are broadly consistent with our expectation and with previous studies of the xillver model (García and Kallman 2010; Dauser et al. 2013): the gas temperature and the strength of the radiation hump show a clear inverse correlation with $\cos\theta_{\mathrm{inc}}$.
@@
III.5 Emergent intensity along all angle bins
Depending on the choice of angular binning, the emergent intensity can be obtained for different emission angles (as was first extensively investigated by García et al. 2014). In the special case where $\mu_{\mathrm{inc}}=\mu_{\mathrm{emis}}$, the emergent radiation field is given by
@@ III.5 Emergent intensity along all angle bins
As shown in Figure 12, the radiation intensity decreases as $\mu$ increases (in agreement with Fig 2 of García et al. 2014) This behavior is primarily driven by the effective optical depth, defined as
@@
Photons escaping at grazing angles (low $\mu$) traverse a longer effective path through the medium. Consequently, spectral features originating from deeper layers, such as emission lines from heavy elements (e.g., Fe), suffer stronger attenuation compared to those produced near the surface (e.g., C, O). For hard X-ray photons, the extended path length increases the probability of scattering, rendering the Compton hump highly sensitive to minor variations in optical depth. This effect, well established in previous works (Lightman et al. 1981; Lightman and White 1988; Magdziarz and Zdziarski 1995), is critical for interpreting the continuum.
The angular dependence of the emissivity plays a vital role in determining the observed spectrum (e.g., García et al. 2014; Liu et al. 2025). Recently, Huang et al. (2025b) highlighted the impact of emission angles on relativistic spectra shaped by strong gravity. In Appendix C, we present relativistic spectra calculated using the formalism adopted from the relconv, relxill and relxillA model.
@@
III.6 Effects of varying the hydrogen density
Currently, DAO assumes reflection in a constant-density slab, with a default hydrogen density of 1015 cm-3. In AGN or XRB systems, the disk densities are sometimes greater than 1015 cm-3 (Jiang et al. 2019) and sometimes even higher than 1020 cm-3 (Liu et al. 2023b). García et al. (2016) and Ding et al. (2024) showed that the emergent reflection spectrum depends sensitively on the assumed disk density. However, the standard XSTAR atomic table (ADTB) only supports hydrogen densities lower than 1018 cm-3; a specific atomic database is needed for higher hydrogen densities (Kallman et al. 2021). DAO will be updated to consider higher densities once the atomic database is available.
In Figure 13, we compare reflection spectra for hydrogen densities of 1015, 1016, and 1017 cm-3 under nthcomp illumination. Other parameters are set at their default values. The results are consistent with the findings of García et al. (2016): for a given ionization parameter, the gas temperature increases with hydrogen density, and the continuum emission in the low energy band displays clear differences among these spectra.
To further study the density effects in the soft X-ray band and temperature profile, we perform additional calculations with a blackbody incident spectrum of temperature 1 keV, which allows us to focus on the soft X-ray band. As shown in Figure 14, the Compton hump is weaker under this type of illumination than for a power-law incident spectrum, and the high-energy parts of the spectra become nearly identical across different densities. However, noticeable differences persist in the low-energy continuum emission. The temperature profiles show obvious differences at large optical depths. Our findings agree with those of García et al. (2022), who introduced the xillverNS model that assumes a blackbody function as the illuminating spectrum.